OpenStudy (anonymous):

Help with a Simultaneous Equation with 3 Variables...

OpenStudy (anonymous):

ok go shoot it here

OpenStudy (anonymous):

\[y+2z-\lambda yz=0\] \[x+2z-\lambda xz=0\] \[2x+2y-\lambda xy=0\] The variables x, y and z should be equated in terms of \[\lambda\]

OpenStudy (anonymous):

Why not give the initial question again and see what the person thinks also

OpenStudy (anonymous):

Let's just look at this as a different question. This is the part I'm stuck on so I've posted this up specifically, so I don't get answers which I don't require.

OpenStudy (anonymous):

ok sorry , fair enuff

OpenStudy (anonymous):

Not having a dig at you or anything, just that I'm stuck on this part, and can't seem to work it out! lol!